OpenStudy (anonymous):

Write the equation of the hyperbola. Foci: (-9,0), (9,0) Vertices: (-4,0), (4,0)

OpenStudy (anonymous):

Hyperbola is a line in which the distance from each focus subtracted gets some constant. \[ \sqrt{(x-9)^2+(y-0)^2} - \sqrt{(x-(-9))^2+(y-0)^2} = c \] We can can find \(c\) by looking at the focus and vertices, since the vertices are on the hyperbola. \[ c = ||-9-4| - |9-4|| = |15 - 5| = 10 \]

OpenStudy (anonymous):

by line, I meant curve

OpenStudy (anonymous):

There is some other equation for it too... something like: \[ \frac{x^2}{a^2}-\frac{y^2}{b^2} = 1 \]But that requires you to know what \(a\) and \(b\) are. I think \(a\) is like the distance of the vertices and \(a/b\) is the slope of the locus... but I am not completely sure.

OpenStudy (anonymous):

@wio thank you so much, what would the final equation be in that case?

OpenStudy (anonymous):

I dunno, I'd have to really try hard at it.